# Multitype entanglement dynamics induced by exceptional points

### v1

@@ VI The calculation for the system with dipolar interaction
We now investigate the EP properties of the system where the two qubits are coupled with dipolar coupling. By numerically solving the matrix of the Hamiltonian Eq. (6), we can obtain the eigenvalues of the system, as shown in Figure 7. In the weak coupling regime (i.e., $g\ll\gamma$) for $g/\gamma=0.006$, there exists an EP2 at $\Omega/\gamma=0.283$ [Fig. 7(a) and  7(b)]. And the EP2 appears at $\Omega/\gamma=0.47$ in the strong coupling regime (i.e., $g\sim\gamma$) for $g/\gamma=0.5$ [Fig. 7(c) and  7(d)].
The $\mathcal{P}\mathcal{T}$-symmetry-broken phase is in the region of $\Omega<{{\Omega}_{\text{EP4}^{\prime}}}$ while the new phase is in the region of ${{\Omega}_{\text{EP4}^{\prime}}}<\Omega<{{\Omega}_{\text{EP2}}}$. It is worth to note that the new phase belongs to the $\mathcal{P}\mathcal{T}$-symmetry-broken phase. Different from the case of Ising type interaction, there is only a single EP2 in the strong coupling regime for the case of dipolar interaction.
@@
For the weak coupling regime, we can find that different entanglement dynamics on the two sides near the original EP4 and EP2 can be observed simultaneously, as shown in Fig. 8(a)-Fig. 8(c). These phenomena show that the different types of entanglement behaviors can still be induced by tuning the order of EP. In the strong coupling regime, the EP2-engineered entanglement transition is still existing [see Fig. 8(d) and  8(e)], but the influence of original EP4 on entanglement behavior is disappeared, as shown in Fig. 8(f). Similar with the results in the case of Ising type interaction, the type I of entanglement behavior only takes place in the weak coupling regime.
Figure 8: For the case of dipolar interaction, the concurrence evolution in the $\mathcal{P}\mathcal{T}$-symmetry-broken phase (a), the new phase (b) and $\mathcal{P}\mathcal{T}$-symmetric phase (c) for weak coupling strength $\xi/\gamma=0.006$. The virtual EP4 induce the exceptional entanglement phenomena on the two sides of this point. In the strong coupling regime for $\xi/\gamma=0.5$, the EP2 induces different entanglement behavior on the two sides of ${{\Omega}_{\text{EP2}}}/\gamma=0.47$, i.e., in the $\mathcal{P}\mathcal{T}$-symmetry-broken phase (d) and $\mathcal{P}\mathcal{T}$-symmetric phase (e). (f) The entanglement dynamics for the driving amplitude $\Omega/\gamma=0.24,0.26$. The initial state is $\left|\psi(0)\right\rangle=\left|aa\right\rangle$.
VII Conclusions
In conclusion, we have discovered multitype entanglement dynamics can be induced by engineering different orders of EP in the system with two coupled NH qubits. It can be found that different entanglement dynamics on the two sides near the original EP4 and EP2 can be observed simultaneously in the weak coupling regime. Moreover, the original EP4-engineered entanglement dynamics only exists in the weak coupling regime and will disappear in the strong coupling regime due to the absence of EP4. We have also demonstrated that an EP-enabled entanglement transition in the system without the driving amplitude in the case of Ising type interaction. Our work provides a protocol for exploring the connection between the entanglement and higher-order EPs, offering opportunities for designing quantum devices by engineering EPs.
@@ Appendix A The concurrence for the two qubits without the driving amplitude
With the increse of the Ising interaction strength $\left|\xi\right|$ until reaching the EP, the energy gap vanishes and both eigenstates approximately converge to the same maximally entangled state as
@@ Appendix B Example for the experimental implementation of Ising type interaction
In our system, we consider the hybrid spin-mechanical system, as shown in Fig. B1(a), where the separated two NV centers are magnetically coupled to the same mechanical motion of a cantilever with dimensions $\left(l,w,t\right)$ via the sharp magnet tip attached to it Pan et al. (2023); Li et al. (2016); P.-B. Li, Y. Zhou, W.-B. Gao, and F. Nori (2020). By applying the cantilever to a periodic drive that modulates its spring constant Rugar and Grütter (1991), it’s possible to amplify the zero-point fluctuations of the mechanical motion. This phenomenon can be experimentally achieved by situating an electrode close to the lower surface of the cantilever and applying an adjustable, time-varying voltage to this electrode Arcizet et al. (2011); Rabl et al. (2010). The electrostatic force gradient stemming from the electrode induces alterations in the spring constant. For a single NV center, the ground-state energy level structure is illustrated in Fig. B1(b). The ground triplet states are $\left|{{m}_{s}}=0,\pm 1\right\rangle$. We applied a homogeneous static magnetic field ${{B}_{static}}$ to remove the degenerate states $\left|{{m}_{s}}=\pm 1\right\rangle$ with the Zeeman splitting $\delta=2{{g}_{e}}{{\mu}_{B}}{{B}_{static}}$, where ${{g}_{e}}\simeq 2$ and ${{\mu}_{B}}=14MHz/mT$ are the $\text{NV}^{\prime}\text{s}$ Lande factor and Bohr magneton, respectively.
Figure B1: Two magnet tips are placed at the end of the silicon cantilever. The spring constant of the cantilever is modified by the electric field from the capacitor plate. Two microwave fields polarized in the x direction (not shown in the picture) are applied to drive the NV centers between the state $\left|{{m}_{s}}=0\right\rangle$ and the states $\left|{{m}_{s}}=\pm 1\right\rangle$. (b) Level diagram of the driven NV center electronic ground state $\left|{{m}_{s}}=0,\pm 1\right\rangle$.
@@
The Hamiltonian for describing the magnetic interaction between the NV spin and the cantilever vibrating mode can be written as ${{H}_{\operatorname{int}}}\equiv{{g}_{e}}{{\mu}_{B}}{{G}_{m}}z{{S}_{z}}$, with ${{G}_{m}}$ the magnetic field gradient. We switch to the dressed state basis $\left|b\right\rangle=1/\sqrt{2}\left(\left|+1\right\rangle-\left|-1\right\rangle\right)$, $\left|g\right\rangle=\cos\theta\left|0\right\rangle-\sin\theta\left|h\right\rangle$, $\left|a\right\rangle=\cos\theta\left|h\right\rangle+\sin\theta\left|0\right\rangle$, with $\left|h\right\rangle=1/\sqrt{2}\left(\left|+1\right\rangle+\left|-1\right\rangle\right)$ and $\tan(2\theta)=-\sqrt{2}\Omega/\Delta$. The lowest energy level $\left|g\right\rangle$ is a stable ground state, which can be treated as an effective continuum outside of the submanifold $\left|a\right\rangle$ and $\left|b\right\rangle$. Furthermore, we assume the transition frequency between the dressed states $\left|a\right\rangle$ and $\left|b\right\rangle$ is almost same as the oscillator frequency, i.e., ${{\omega}_{ab}}\sim{{\omega}_{m}}$. This system can be described as
@@
$\quad+\sum\limits_{i=1,2}\left[{{g}_{i}}\left({{a}^{\dagger}}\sigma_{i}^{-}+a\sigma_{i}^{\dagger}\right)+\frac{{{\delta}_{dg}}}{2}\sigma_{z}^{i}\right]$ (11)
where the coefficients are ${{\delta}_{m}}={{\omega}_{m}}-{{\omega}_{p}}$, ${{\delta}_{ab}}={{\omega}_{ab}}-{{\omega}_{p}}$, $g={{g}_{e}}{{\mu}_{B}}{{G}_{m}}{{z}_{zpf}}\sin\theta$, $\sigma_{z}^{i}\equiv{{\left|a\right\rangle}_{i}}\left\langle a\right|-{{\left|b\right\rangle}_{i}}\left\langle b\right|$, $\sigma_{i}^{+}\equiv{{\left|a\right\rangle}_{i}}\left\langle b\right|$ and $\sigma_{i}^{-}\equiv{{\left|b\right\rangle}_{i}}\left\langle a\right|$. Therefore, it is obvious that the single NV can be seen as the two-level emitter. By applying another classical field to drive the two spins (with amplitude $\Omega$), the driving Hamiltonian can be obtained as ${{H}_{dri}}=\Omega(\sigma_{x}^{1}\text{+}\sigma_{x}^{2})$. Thus, the total Hamiltonian of our system can be rewritten as
@@
$\quad+\left.\frac{{{\delta}_{dg}}}{2}\sigma_{z}^{i}\right]+{{H}_{dri}}$
@@
Considering ${{\delta}_{dg}}=0$ and ${{g}_{eff,1}}={{g}_{eff,2}}={{g}_{eff}}$, the Rabi Hamiltonian $H_{Rabi}^{S}$ can be reduced using the Schrieffer-Wolff transformation $H_{Rabi}^{eff}={{e}^{S}}H_{Rabi}^{S}{{e}^{-S}}$ where $S=\chi(a_{s}^{\dagger}-{{a}_{s}})(\sigma_{x}^{1}+\sigma_{x}^{2})$ and $\chi={{g}_{eff}}/{{\Delta}_{m}}$. Noted that the parameter $\chi$ is much smaller than one, indicating that it satisfies the Lamb-Dicke condition $\chi\ll 1$. The effective Hamiltonian is given by
$H_{RT}^{eff}={{\Delta}_{m}}a_{s}^{\dagger}{{a}_{s}}-\xi{{\left(\sigma_{x}^{1}+\sigma_{x}^{2}\right)}^{2}}+\Omega(\sigma_{x}^{1}+\sigma_{x}^{2}).$ (14)
where $\xi=g_{eff}^{2}/{{\Delta}_{m}}$. Owing to the operator ${{a}_{s}}$ is decoupled from the two-level emitter, we can only retain the second and third terms in Eq. (13). Especially, the second term is the typical Ising type interaction Hamiltonian ${{H}_{Ising}}=-\xi{{\left(\sigma_{x}^{1}+\sigma_{x}^{2}\right)}^{2}}$. Therefore, we rewritten the Hamiltonian Eq. (14) as
${H}_{eff}=-\xi{{(\sigma_{x}^{1}+\sigma_{x}^{2})}^{2}}+\Omega(\sigma_{x}^{1}+\sigma_{x}^{2}).$ (15)
@@
Figure B2: The concurrence evolution of the two spins system by respectively solving the two subequations of Eq. (16). The initial state of the system is $\left|0\right\rangle\left|aa\right\rangle$ and $\left|aa\right\rangle$ for Eq. (16)(a) and Eq. (16)(b), respectively. The left parameters are ${{\delta}_{dg}}=0$, $\kappa_{a}^{S}=\gamma_{NV}^{i}=\gamma$, $\Omega/\gamma=0.3$, $\xi/\gamma=0.0006$, ${{\Delta}_{m}}/\gamma=40$ and ${{g}_{eff}}=\sqrt{\xi{{\Delta}_{m}}}$.
@@ Appendix C Non-Hermitian perturbation theory in Ising type interaction model
The non-Hermiticity of $H$ is the mean reason for the non-orthogonality, in order to better address the importance of non-Hermiticity, we extend the state basis to include states in the dual Hilbert space ${{H}^{*}}$. The Hamiltonian in the dual space ${{H}^{*}}$ is just the Hermitian conjugate of ${{H}_{j}}$ in Eq. (17). i.e., ${{\overline{H}}_{j}}=H_{j}^{\dagger}$, with the eigenvalues ${{\overline{\lambda}}_{\pm}}=\frac{1}{4}\left(i\gamma-4\xi\pm\sqrt{16{{\Omega}^{2}}-{{\gamma}^{2}}}\right)$ and the corresponding eigenstates
@@
with the corresponding eigenstates are
@@
For the eigenvalues and eigenstates of ${{\overline{H}}_{0}}$ in the dual space are given by,

### v2

@@ VII The calculation for the system with dipolar interaction
We now investigate the relationship with the EP properties and entanglement behavior of the system, where the two qubits are coupled with dipolar coupling. By numerically solving the matrix of the Hamiltonian  (10), we can obtain the eigenvalues of the system, as shown in Figure 7. In the weak coupling regime (i.e., $g\ll\gamma$) for $g/\gamma=0.006$, there exists an EP2 at $\Omega/\gamma=0.283$ [Figs. 7(a) and  7(c)]. The region of PI is in $\Omega<{{\Omega}_{\text{EP4}^{\prime}}}$ while the mixed phase is in the region of ${{\Omega}_{\text{EP4}^{\prime}}}<\Omega<{{\Omega}_{\text{EP2}}}$. Similarly, both the PI and mixed phase belong to the PTB. Only a single EP2 appears at $\Omega/\gamma=0.47$ in the strong coupling regime (i.e., $g\sim\gamma$) for $g/\gamma=0.5$ [Fig. 7(b) and  7(d)].
@@
In the weak coupling regime, we can still observe the different entanglement dynamics change at the EP4’ and EP2, similarly with the model of Ising type interaction system, as shown in Fig. 8(a)-Fig. 8(c). Correspondingly, the characteristic of the entanglement dynamics also depends on the feature of the eigenvalues.
In the strong coupling regime, the EP2 induces the spectrum transition of the PTS and mixed phase, leading to the type II [Fig. 8(e)] and III [Fig. 8(d)] entanglement dynamics change. However, the influence of original EP4 on entanglement behavior is disappears, indicating that only type III entanglement dynamics can be observed, as shown in Fig. 8(f). Similarly, as in the case of Ising type interaction, type I entanglement dynamics only occur in the weak coupling regime
Figure 8: For the case of dipolar interaction, the concurrence evolution in PI (a), the mixed phase (b) and PTS (c) for weak coupling strength $\xi/\gamma=0.006$. The EP4’ and EP2 induce the different entanglement dynamics changes on the two sides of these points. In the strong coupling regime for $\xi/\gamma=0.5$, the EP2 induces the spectrum transition of the PTS and mixed phase, leading to the type II (e) and III (d) entanglement dynamics change. The entanglement dynamics for the driving amplitude $\Omega/\gamma=0.24,0.26$ (f). All the initial states are $\left|\psi(0)\right\rangle=\left|aa\right\rangle$.
VIII Conclusions
In conclusion, we have found that entanglement dynamics change can be induced by the EPs in a system of two coupled NH qubits. Divided by EP2 and the original EP4, three distinct types of entanglement dynamics can be observed. By analyzing the real and imaginary parts of the eigenvalues of the system, we demonstrate that the characteristics of the eigenvalues determine the types of the entanglement dynamics. The original EP4 induces the spectrum transition of the PI and mixed phase, leading to the type I and III entanglement dynamics change in the weak coupling regime. Similarly, the EP2 induces the spectrum transition of the PTS and mixed phase, leading to the type II and III entanglement dynamics change in the both weak and strong coupling regimes. Furthermore, as the eigenstates are entangled states, we further confirm that the concurrence associated with the eigenstates can be treated as another manifestation of the entanglement dynamics change. Moreover, we have also proved that an EP-enabled entanglement changes in the system without the driving amplitude in the case of Ising-type interaction. Our work provides a protocol for exploring the connection between the entanglement and higher-order EPs, offering opportunities for designing quantum devices by engineering EPs.
@@ Appendix A The concurrence for the two qubits without the driving amplitude
With the increase of the Ising interaction strength $\left|\xi\right|$ until reaching the EP, the energy gap vanishes and both eigenstates approximately converge to the same maximally entangled state as
@@
In the main text, we can observe the different entanglement dynamics around the EP. Physically speaking, these two different behaviors induced by the EP can be understood as a consequence of the competition between the Ising-type interaction and incoherent dissipation.
@@ Appendix B Example for the experimental implementation of Ising type interaction
In our system, we consider the hybrid spin-mechanical system, as shown in Fig. B1(a), where the separated two NV centers are magnetically coupled to the same mechanical motion of a cantilever with dimensions $\left(l,w,t\right)$ via the sharp magnet tip attached to it X.-F. Pan, X.-L. Hei, X.-L. Dong, J.-Q. Chen, C.-P. Shen, H. Ali, and P.-B. Li (2023); P. B. Li, Z.-L. Xiang, P. Rabl, and F. Nori (2016); P. B. Li, Y. Zhou, W. B. Gao, and F. Nori (2020). By applying the cantilever to a periodic drive that modulates its spring constant Rugar and Grutter (1991), it is possible to amplify the zero-point fluctuations of the mechanical motion. This phenomenon can be experimentally achieved by situating an electrode close to the lower surface of the cantilever and applying an adjustable, time-varying voltage to this electrode O. Arcizet, V. Jacques, A. Siria, P. Poncharal, P. Vincent, and S. Seidelin (2011); P. Rabl, S. J. Kolkowitz, F. H. L. Koppens, J. G. E. Harris, P. Zoller, and M. D. Lukin (2010). The electrostatic force gradient stemming from the electrode induces alterations in the spring constant. For a single NV center, the ground-state energy level structure is illustrated in Fig. B1(b). The ground triplet states are $\left|{{m}_{s}}=0,\pm 1\right\rangle$. We applied a homogeneous static magnetic field ${{B}_{static}}$ to remove the degenerate states $\left|{{m}_{s}}=\pm 1\right\rangle$ with the Zeeman splitting $\delta=2{{g}_{e}}{{\mu}_{B}}{{B}_{static}}$, where ${{g}_{e}}\simeq 2$ and ${{\mu}_{B}}=14MHz/mT$ are the $\text{NV}^{\prime}\text{s}$ Lande factor and Bohr magneton, respectively.
Figure B1: (a) Two magnet tips are placed at the end of the silicon cantilever. The spring constant of the cantilever is modified by the electric field from the capacitor plate. Two microwave fields polarized in the x direction (not shown in the picture) are applied to drive the NV centers between the state $\left|{{m}_{s}}=0\right\rangle$ and the states $\left|{{m}_{s}}=\pm 1\right\rangle$. (b) Level diagram of the driven NV center electronic ground state $\left|{{m}_{s}}=0,\pm 1\right\rangle$.
@@
The Hamiltonian for describing the magnetic interaction between the NV spin and the cantilever vibrating mode can be written as ${{H}_{\operatorname{int}}}\equiv{{g}_{e}}{{\mu}_{B}}{{G}_{m}}z{{S}_{z}}$, with ${{G}_{m}}$ the magnetic field gradient. We assume that the two NV centers symmetrically placed on either side of the micromagnet along the direction of magnetization, and only consider the coupling between the oscillation in the $z$ direction and the NV center. Thus, the interaction between the NV centers can be limited in the $x$ direction X.-F. Pan, X.-L. Hei, X.-L. Dong, J.-Q. Chen, C.-P. Shen, H. Ali, and P.-B. Li (2023). We switch to the dressed state basis $\left|b\right\rangle=1/\sqrt{2}\left(\left|+1\right\rangle-\left|-1\right\rangle\right)$, $\left|g\right\rangle=\cos\theta\left|0\right\rangle-\sin\theta\left|h\right\rangle$, $\left|a\right\rangle=\cos\theta\left|h\right\rangle+\sin\theta\left|0\right\rangle$, with $\left|h\right\rangle=1/\sqrt{2}\left(\left|+1\right\rangle+\left|-1\right\rangle\right)$ and $\tan(2\theta)=-\sqrt{2}\Omega/\Delta$. The lowest energy level $\left|g\right\rangle$ is a stable ground state, which can be treated as an effective continuum outside of the submanifolds $\left|a\right\rangle$ and $\left|b\right\rangle$. Furthermore, we assume the transition frequency between the dressed states $\left|a\right\rangle$ and $\left|b\right\rangle$ is almost the same as the oscillator frequency, i.e., ${{\omega}_{ab}}\sim{{\omega}_{m}}$. This system can be described as
@@
$\quad+\sum\limits_{i=1,2}\left[{{g}_{i}}\left({{a}^{\dagger}}\sigma_{i}^{-}+a\sigma_{i}^{\dagger}\right)+\frac{{{\delta}_{dg}}}{2}\sigma_{i}^{z}\right]$ (15)
where the coefficients are ${{\delta}_{m}}={{\omega}_{m}}-{{\omega}_{p}}$, ${{\delta}_{ab}}={{\omega}_{ab}}-{{\omega}_{p}}$, $g={{g}_{e}}{{\mu}_{B}}{{G}_{m}}{{z}_{zpf}}\sin\theta$, $\sigma_{i}^{z}\equiv{{\left|a\right\rangle}_{i}}\left\langle a\right|-{{\left|b\right\rangle}_{i}}\left\langle b\right|$, $\sigma_{i}^{+}\equiv{{\left|a\right\rangle}_{i}}\left\langle b\right|$ and $\sigma_{i}^{-}\equiv{{\left|b\right\rangle}_{i}}\left\langle a\right|$. Therefore, it is obvious that the single NV can be seen as the two-level emitter. By applying another classical field to drive the two spins (with amplitude $\Omega$), the driving Hamiltonian can be obtained as ${{H}_{dri}}=\Omega(\sigma_{1}^{x}\text{+}\sigma_{2}^{x})$. Thus, the total Hamiltonian of our system can be rewritten as
@@
$\quad+\left.\frac{{{\delta}_{dg}}}{2}\sigma_{i}^{z}\right]+{{H}_{dri}}$
@@
Considering ${{\delta}_{dg}}=0$ and ${{g}_{eff,1}}={{g}_{eff,2}}={{g}_{eff}}$, the Rabi Hamiltonian $H_{Rabi}^{S}$ can be reduced using the Schrieffer-Wolff transformation $H_{Rabi}^{eff}={{e}^{S}}H_{Rabi}^{S}{{e}^{-S}}$ where $S=\chi(a_{s}^{\dagger}-{{a}_{s}})(\sigma_{1}^{x}+\sigma_{2}^{x})$ and $\chi=\frac{{{g}_{eff}}}{{{\Delta}_{m}}}$. Noted that the parameter $\chi$ is much smaller than one, indicating that it satisfies the Lamb-Dicke condition $\chi\ll 1$. The effective Hamiltonian is given by
$H_{RT}^{eff}={{\Delta}_{m}}a_{s}^{\dagger}{{a}_{s}}-\xi{{\left(\sigma_{1}^{x}+\sigma_{2}^{x}\right)}^{2}}+\Omega(\sigma_{1}^{x}\text{+}\sigma_{2}^{x}).$ (18)
where $\xi=\frac{g_{eff}^{2}}{{{\Delta}_{m}}}$. Owing to the operator ${{a}_{s}}$ is decoupled from the two-level emitter, we can only retain the second and third terms in Eq. (17). Especially, the second term is the typical Ising type interaction Hamiltonian ${{H}_{Ising}}=-\xi{{\left(\sigma_{1}^{x}+\sigma_{2}^{x}\right)}^{2}}$. Therefore, we rewrite the Hamiltonian (18) as
${H}_{eff}=-\xi{{(\sigma_{1}^{x}\text{+}\sigma_{2}^{x})}^{2}}+\Omega(\sigma_{1}^{x}\text{+}\sigma_{2}^{x}).$ (19)
@@
Figure B2: The concurrence evolution of the two spins system by respectively solving the two subequations of Eq. (20). The initial state of the system is $\left|0\right\rangle\left|aa\right\rangle$ and $\left|aa\right\rangle$ for Eq. (20)(a) and Eq. (20)(b), respectively. And the coefficients are ${{\delta}_{dg}}=0$, $\kappa_{a}^{S}=\gamma_{NV}^{i}=\gamma$, $\Omega/\gamma=0.3$, $\xi/\gamma=0.0006$, ${{\Delta}_{m}}/\gamma=40$ and ${{g}_{eff}}=\sqrt{\xi{{\Delta}_{m}}}$.
@@ Appendix C Non-Hermitian perturbation theory in Ising type interaction model
The non Hermiticity of $H$ is the main reason for the non-orthogonality, in order to better address the importance of non-Hermiticity, we extend the state basis to include states in the dual Hilbert space ${{H}^{*}}$. The Hamiltonian in the dual space ${{H}^{*}}$ is just the Hermitian conjugate of ${{H}_{j}}$ in Eq. (21). i.e., ${{\overline{H}}_{j}}=H_{j}^{\dagger}$, with the eigenvalues ${{\overline{\lambda}}_{\pm}}=\frac{1}{4}\left(i\gamma-4\xi\pm\sqrt{16{{\Omega}^{2}}-{{\gamma}^{2}}}\right)$ and the corresponding eigenstates
@@
with the corresponding eigenstates
@@
The eigenvalues and eigenstates of ${{\overline{H}}_{0}}$ in the dual space are given by,